AAVSOnet processing update
We were waiting for new flats for W28, W30, K28, BSM, which were taken on 2011-01-10 (110110). The images from 110104, 110110, 110111 have now been processed and are available; new data from last night (110113) will be available shortly. K28 processing will resume this weekend and will be caught up by Monday. SRO50 should be up to date.